HETAV PATEL

Grade 11,

The enthalpy of fusion of ice is 6kJ/mol. The heat capacity of water is 4.2j/g/K.What is the smallest amount of ice cubes at 0degrees,each containing i mole of water,that are needed to cool 500g of liquid water from 20 dgreess to 0degrees Ans:7.Give me the complete solution

The enthalpy of fusion of ice is 6kJ/mol. The heat capacity of water is 4.2j/g/K.What is the smallest amount of ice cubes at 0degrees,each containing i mole of water,that are needed to cool 500g of liquid water from 20 dgreess to 0degrees


Ans:7.Give me the complete solution

Grade:12

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

heat rejected by water = m*c*delT = 500*4.2*20

let m be the number of cubes

heat taken by cubes = m*6000

heat rejected = heat raken

=> m*6000 = 500*4.2*20

=> m = 7